OpenStudy (safyousuff97):

Part A: Max rented a motorbike at $465 for 5 days. If he rents the same motorbike for a week, he has to pay a total rent of $625. Write an equation in the standard form to represent the total rent (y) that Max has to pay for renting the motorbike for x days. (4 points) Part B: Write the equation obtained in Part A using function notation. (2 points) Part C: Describe the steps to graph the equation obtained above on the coordinate axes. Mention the labels on the axes and the intervals. (4 points)

OpenStudy (safyousuff97):

@AaronAndyson

OpenStudy (aaronandyson):

part A: y=65+80(x) plug in 5. 65+80(5) y=65+400. y=465. when he rents the bike for five days, he pays 465. so if he rents it for a week he would be renting it for7 days (x)days. so y=65+80(7) 65+560. y=625.

OpenStudy (aaronandyson):

Now find the slope.

OpenStudy (aaronandyson):

A.625-465=$160 he has to pay for 2 days=160 $ rent paid for 1 day=260/2=$80 amount paid for 5days=8*5=$400 But he paid $465 Difference 465-400=$65 so every one has to pay $65+rent for the no. of days total rent for x days y=65+80 x (dollars) B. f(x)=65+80 x

OpenStudy (safyousuff97):

how do i describe the steps ? for part c

OpenStudy (aaronandyson):

Plot the following points (4,465) (7,625)